# A Nonparametric Bayesian Solution of the Empirical Stochastic Inverse Problem

### v1

@@ Abstract
The stochastic inverse problem is a key ingredient in making inferences, predictions, and decisions for complex science and engineering systems. We formulate and analyze a nonparametric Bayesian solution for the stochastic inverse problem. Key properties of the solution are proved and the convergence and error of a computational solution obtained by random sampling is analyzed. Several applications illustrate the results.
@@ 1 Introduction
The behaviour of a complex system is governed by physical processes and laws, while the behaviour of any specific example of the system depends on its unique characteristics, or parameters, e.g., geometry, material properties, or rates of interaction. For instance, many materials respond to heat stress in similar ways, but the response of a particular object depends on factors such as its composition, thermal diffusivity, and shape.
@@
In many real-world settings, parameters determining the behavior of a complex system cannot be measured directly. Instead, experiments typically produce observations of system behavior. To make sense of these observations, scientists and engineers rely on computer models that encode the mathematical representations of the governing physical laws to link parameters to observations. For example, the heat equation, a well-validated partial differential equation, describes how an object responds to a heat source. Measuring the object’s temperature at a given time and location corresponds to evaluating the solution of this model, and from such measurements one can infer information about parameters related to the object’s material composition such as thermal conductivity.
@@
This gives rise to the general inverse problem where investigators infer the unknown parameters using a process model and observations. Solving this problem plays an important role in many applications, e.g., forecasting epidemics, designing new materials, assessing radiation damage in nuclear fuel rods, predicting climate and weather, detection of black holes, modeling subsurface contamination in aquifers, and estimating storm surge from hurricanes. There are many approaches to this inverse problem; see Sec. 1.4 for more details. We take a statistical approach in which the inferential object of interest is the distribution of parameters. Obtaining a distribution is a key ingredient for predicting system behavior.
1.1 The empirical stochastic inverse problem for random experiments
@@
We begin by introducing the inverse problem and giving an overview of the main results. We assume that the parameters determining system behavior take values in a set $\Lambda\subset\mathbb{R}^{n}$, $n\geq 1$. We consider the common situation in which it is not possible to observe the parameters directly. Instead, we assume that it is possible to observe data that corresponds to a function of the parameters. Specifically, we observe values of a vector or scalar function $Q(\lambda)$ for any $\lambda\in\Lambda$, where $Q:\Lambda\to\mathcal{D}$ with range $\mathcal{D}=Q(\Lambda)\subset\mathbb{R}^{m}$ for $m\leq n$. $Q$ is called the quantities (quantity) of interest (QoI).
@@
In many cases, defining the QoI is key to experimental design. As with the heat equation, $Q(\lambda)=q(Y(\lambda))$ is typically obtained by applying a observation map $q$ to the solution $Y$ of a process model $\mathcal{M}(Y,\lambda)=0$ that encapsulates physical properties, e.g., conservation of mass and energy, where $Y$ is an implicit function of the physical parameters $\lambda\in\Lambda$. An important issue is that $Q$ is generally not 111-1, which reflects the loss of information in experimental observations. In this case, multiple physical instances, i.e., values of $\lambda$, yield the same set of experimental data. It turns out that this is directly related to the introduction of conditional probability.
@@
For any $q\in\mathcal{D}$, the generalized contour $Q^{-1}(q)$ gives all possible input parameter values $\lambda$ that result in the same output. It is the generalization of a contour curve for a scalar map on two dimensions. Because $\mathcal{D}$ is the range of $Q$, we obtain the decomposition $\Lambda=\bigcup_{q\in\mathcal{D}}Q^{-1}(q)$, where $Q^{-1}(q_{1})\cap Q^{-1}(q_{2})=\emptyset$ when $q_{1}\neq q_{2}$. Below, we investigate the properties of generalized contours that are important for the practical use of conditional probability.
We now have the ingredients for the classic model of a random experiment. Given data $\{q_{i}\}_{i=1}^{K}$ for $K$ trials of the experiment, we assume that there is a collection of $K$ parameter values $\{\lambda_{i}\}_{i=1}^{K}$ that are distributed according to $P^{\mathrm{d}}_{\Lambda}$ such that $q_{i}=Q(\lambda_{i})$ for $1\leq i\leq K$. The collection of data $\{q_{i}\}_{i=1}^{K}$ is distributed according to $P_{\mathcal{D}}$. We define the stochastic inverse problem in the situation that $P^{\mathrm{d}}_{\Lambda}$ is unknown. In this context, $P^{\mathrm{d}}_{\Lambda}$ is called the data generating distribution.
@@
The eSIP has multiple solutions when $Q$ is not 111-1. We adopt a Bayesian approach to determine a unique solution of the eSIP conditioned on a prior distribution that encapsulates prior information about the physical system.
@@
We return to consider the eSIP when developing the computational solution methodology in Sec. 3.
@@ 1.2 An overview of the main results
Assume that $P_{\mathcal{D}}$ has density $\rho_{\mathcal{D}}$ with respect to Lebesgue measure. Under general conditions, for each prior distribution $P_{\mathrm{p}}$ on $(\Lambda,\mathcal{B}_{\Lambda})$ with density $\rho_{\mathrm{p}}$ there is a unique posterior distribution ${P}_{\Lambda}$ that solves the SIP and there is an explicit formula for its density ${\rho}_{\Lambda}$ determined from $\rho_{\mathcal{D}}$ and $\rho_{\mathrm{p}}$. This is the nonparametric analog of the result typically deployed for classic Bayesian statistics.
@@
Under general conditions, the density ${\rho}_{\Lambda}$ of the Bayesian solution of the SIP is continuous a.e.
@@
Under general conditions, the Bayesian solution of the SIP corresponding to the uniform prior has maximum entropy.
@@
Returning to the eSIP, we construct a numerical estimate $\hat{P}_{\Lambda}\approx P_{\Lambda}$ based on a form of importance sampling. Under general conditions, the estimator $\hat{P}_{\Lambda}$ is asymptotically unbiased and strongly consistent and we provide asymptotic estimates of its variance and error.
Broadly speaking, the first result says that even for complex computer models, there exists a unique posterior distribution to the SIP with an explicit formula for the density. The second result implies that the posterior distribution is well-behaved in practical settings. The third main result says that under general conditions that the uniform prior is the least informative choice. Finally the fourth main result is that there is a robust, reliable, and practical numerical solution methodology.
@@
We compute the posterior distribution solving the eSIP with the uniform prior using the numerical solution method presented in Sec. 3 with a large number of samples in order to virtually eliminate effects of finite sampling. We sample $16\times 10^{6}$ points in $\Lambda$ randomly from $P^{\mathrm{d}}_{\Lambda}$. We use the synthetic data to build an empirical output distribution on a partition of $\mathcal{D}$ with $100$ cells. By a partition of a measurable set $A$, we mean a finite or countable collection of non-intersecting measurable subsets $\{B_{i}\}_{i}$ such that $A=\bigcup_{i}B_{i}$, except possibly for set of measure 0. We compute an empirical solution of the eSIP using $16\times 10^{6}$ uniformly distributed points in $\Lambda$.
@@
In Fig. 3, we plot heatmaps of the eSIP solutions for $T=.5$ and $T=2$ produced by computing $\{P_{\Lambda}(B_{i})\}_{i=1}^{6400}$ for a partition $\{B_{i}\}_{i=1}^{6400}$ of $\Lambda$ into small cells $B_{i}$ and generating a heatmap from the probabilities. Recalling the contours shown in Fig. 1, the solutions have the characteristic “ridge” shape oriented along contours that reflects the fact that parameter values on the same contour have equal probability due to the choice of the uniform prior. This reflects the indeterminacy induced by the loss of information in the experimental observation. Note that we do not expect to recover the data generating distribution. We compute a posterior solution of the SIP conditioned on the choice of prior.
@@
Figure 3: Left and center: Heatmaps of solutions of the eSIP for the exponential decay model using the uniform prior for $T=.5$ (left) and $T=2$ (center). Right: Overlays of the two heatmaps. Event A is relatively low probability for both solutions, Event B is relatively high probability for the solution for time .5.5 but relatively low probability for the solution for time 22, while Event C is the reverse. Event D is relatively high probability for both solutions.
1.3 Using the solution of the eSIP
@@
In applications of the eSIP, the principal inferential target is information about the probability of combinations of parameters consistent with the model and the observations. The solution of the eSIP is a posterior probability distribution on the sample space $\Lambda$ that is evaluated to estimate probabilities of events. For example in the right hand plot in Fig. 3, we show four events of equal size located in different parts of $\Lambda$. We see that event $A$ has relatively low probability with respect to the eSIP solutions for $T=.5$ and $T=2$, while event $D$ has relatively high probability for both eSIP solutions. Events $B$ and $C$ are relatively high probability for one of the eSIP solutions but not the other.
@@
We emphasize that the solution of the eSIP is a posterior probability distribution, not a point estimate. We can compute statistics of the posterior such as a mean or variance but those are typically of little use.
@@
The other use of the solution of the eSIP is forecasting or prediction, which means computing a probability distribution on a new quantity of interest. For example, we can compute the solution of the eSIP for observations at an early time, then use the solution to estimate the induced distribution on the quantity of interest at a later time. This is related to ensemble forecasting, see [46] for an example of forecasting infections of COVID during a surge using real data.
@@ 1.4 Observations and related work
The data for the eSIP results from conducting trials of the experiment multiple times, with each trial conducted with physical parameters distributed according to an unknown data generating distribution. The eSIP is closely related to ensemble forecasting, e.g., as used in weather prediction [27, 1, 50, 49, 48, 34, 55, 46]. The stochastic variation in output data is the result of random sampling of the physical parameters rather than from observational noise. Thus, the eSIP is significantly different than the common statistics problem in which stochastic data results from noisy evaluation of an experiment conducted for a single set of parameters (a “true” value). Observational error can be added to the eSIP, and more generally, the eSIP can be extended to include random effects, but this increases the complexity and dimension of the inverse problem to be solved [5].
The inferential target of the eSIP is a probability distribution on the physical parameters as opposed to a distribution on parameters in a distributional model in a classic application of Bayesian statistics. The sample space $\Lambda$ commonly includes ranges of values that lead to significantly different behavior in the system, e.g., $\Lambda$ may include bifurcation points. The distribution of the QoI is correspondingly complex. In general, the solution of the eSIP is a probability distribution with a complex heterogeneous structure, motivating the use of a nonparametric approach to computing approximate solutions.
The solution depends on disintegration of measures rather than an application of Bayes theorem [41, 21, 63, 8, 43, 58, 4]. The regular conditional probability resulting from conditioning on the $\sigma$-algebra induced by $Q$ plays a prominent role in the construction of the Bayesian solution of the eSIP. This reflects the fact that the experimental data provides limited information about the physical parameters.
The review paper [6] explains the relationship between the eSIP and the so-called Bayesian Inverse Problem widely employed in engineering [31, 24, 66, 28, 19, 68, 69]. The Bayesian Inverse Problem is a type of stochastic backward error analysis. While it supposedly employs Bayes’ Theorem in some fashion, it is not a statistical problem insofar as it is not generally based on field data nor Bayesian in a classical statistical sense. The solution depends on regularization.
Bayesian computer model calibration is a statistical approach that combines experimental data and computer model output to estimate a ”true” value for the parameters (“calibration parameters”) in the presence of uncertainty [44, 38, 39, 70, 57, 37, 51, 74, 7]. Bayesian calibration typically employs an “emulator” (a statistical surrogate model for the computer model) and often attempts to model the systematic “discrepancy” between the calibrated model emulator and the mean of the physical system. The work in [37], for example, considers an inverse problem similar to the eSIP.
Another form of inverse problem arises when only longitudinal data for a single trial of an experiment is available. Solution methodologies for this kind of inverse problem include 4D-VAR, data assimilation, ensemble Kalman filtering, Bayesian particle filtering, et cetera [1, 33, 2, 23, 3, 20, 30, 30, 71, 26, 61]. All of these approaches adopt (typically unverifiable) model assumptions and solution formulations based on regularization. So, both the problem and solution approaches are much different than the eSIP.
@@
The original development of the SIP used a different form of disintegration and employed a deterministic numerical solution method [9, 11, 12]. This paper introduces a new form of disintegration of the solution that enables derivation of some key properties for the first time as well as analysis of a novel numerical solution method based on random sampling. The advances presented in this paper depend on the developments in the theses [73, 22]. Extensions of the eSIP to other problems are considered in [53, 17, 73, 22]. Consideration of computational aspects of the solution of the eSIP are considered in [18, 15, 16, 65, 76, 77]. Applications of the eSIP can be found in [10, 14, 13, 52, 35, 42, 46]. In traditional inverse problems, “ill-conditioning” and the condition number play important roles. In [65], we develop a theory of condition for the eSIP and explore the consequences for the solution. In [5], we consider the inclusion of random effects in the eSIP, which opens up doors to numerous applications and extensions, including connecting the eSIP to traditional Bayesian statistics.

### v3

# Nonparametric Bayesian Calibration of Computer Models

@@ Abstract
Calibration of computer models is a key step in making inferences, predictions, and decisions for complex science and engineering systems. We formulate and analyze a nonparametric Bayesian methodology for computer model calibration. This paper presents a number of key results including; establishment of a unique nonparametric Bayesian posterior corresponding to a chosen prior with an explicit formula for the corresponding conditional density; a maximum entropy property of the posterior corresponding to the uniform prior; the almost everywhere continuity of the density of the nonparametric posterior; and a comprehensive convergence and asymptotic analysis of an estimator based on a form of importance sampling. We illustrate the problem and results using several examples, including a simple experiment.
@@ 1 Introduction
The behaviour of a complex system is governed by physical processes and laws, while the behaviour of any specific example of the system depends on its unique characteristics, or (physical) parameters, e.g., geometry, material properties, or rates of interaction. For instance, many materials respond to heat stress in similar ways, but the response of a particular object depends on factors such as its composition, thermal diffusivity, and shape.
@@
In many real-world settings, parameters determining the behavior of a complex system cannot be measured directly. Instead, experiments typically produce observations of system behavior. To make sense of these observations, scientists and engineers rely on computer models that encode the mathematical representations of the governing physical laws to link parameters to observations. For example, the heat equation is a well-validated partial differential equation that describes how an object responds to a heat source. Measuring the object’s temperature at a given time and location corresponds to evaluating the solution of this model, and from such measurements one can infer information about parameters related to the object’s material composition such as thermal conductivity. In practice, the solution of a complex differential equation is approximated using a computer model.
@@
This gives rise to the general calibration of inferring the unknown parameters using a computer model and observed data. Solving this problem plays an important role in many applications, e.g., forecasting epidemics, designing new materials, assessing radiation damage in nuclear fuel rods, predicting climate and weather, detection of black holes, modeling subsurface contamination in aquifers, and estimating storm surge from hurricanes. We take a statistical approach to calibration in which the inferential object of interest is the distribution of parameters. Obtaining a distribution on the physical parameters is a key ingredient for predicting system behavior.
1.1 Calibration and the empirical stochastic inverse problem
@@
We begin by formulating calibration of a computer model as a form of inverse problem and giving an overview of the main results for its solution. We assume that the parameters determining system behavior take values in a set $\Lambda\subset\mathbb{R}^{n}$, $n\geq 1$. We consider the common situation in which it is not possible to observe the parameters directly. Instead, we assume that it is possible to observe data that corresponds to a function of the parameters. Specifically, we observe values of a vector or scalar function $Q(\lambda)$ for any $\lambda\in\Lambda$, where $Q:\Lambda\to\mathcal{D}$ with range $\mathcal{D}=Q(\Lambda)\subset\mathbb{R}^{m}$ for $m\leq n$. $Q$ is called the quantities (quantity) of interest (QoI).
@@
In many cases, defining the QoI is key to experimental design. Formally, $Q(\lambda)=q(Y(\lambda))$ is typically obtained by applying a observation map $q$ to the solution $Y$ of a process model $\mathcal{M}(Y,\lambda)=0$ that encapsulates physical properties, e.g., conservation of mass and energy, where $Y$ is an implicit function of the physical parameters $\lambda\in\Lambda$. The solution of the process model is approximated using a computer model. We use $Q$ to refer to both the underlying process model and the related computer model. An important issue is that $Q$ is generally not 111-1, which reflects the loss of information in experimental observations. In this case, multiple physical instances, i.e., values of $\lambda$, yield the same set of experimental data. It turns out that this is directly related to the introduction of conditional probability.
@@
For any $q\in\mathcal{D}$, the generalized contour $Q^{-1}(q)$ gives all possible input parameter values $\lambda$ that result in the same output. It is the generalization of a contour curve for a scalar map on two dimensions. Note that $Q^{-1}(q_{1})\cap Q^{-1}(q_{2})=\emptyset$ when $q_{1}\neq q_{2}$ and $\Lambda=\bigcup_{q\in\mathcal{D}}Q^{-1}(q)$.
We now have the ingredients for the classic model of a random experiment. Given data $\{q_{i}\}_{i=1}^{K}$ for $K$ trials of the experiment, we assume that there is a collection of $K$ parameter values $\{\lambda_{i}\}_{i=1}^{K}$ that are distributed according to $P^{\mathrm{d}}_{\Lambda}$ such that $q_{i}=Q(\lambda_{i})$ for $1\leq i\leq K$. The collection of data $\{q_{i}\}_{i=1}^{K}$ is distributed according to $P_{\mathcal{D}}$. In this context, $P^{\mathrm{d}}_{\Lambda}$ is called the data generating distribution because each sample from $P^{\mathrm{d}}_{\Lambda}$ sets the experimental conditions for a trial resulting in an observed datum.
The calibration problem for the computer model arises in the situation that $P^{\mathrm{d}}_{\Lambda}$ is unknown. In this case, calibration can be framed as an inverse problem.
@@
The eSIP has multiple solutions when $Q$ is not 111-1. We adopt a Bayesian approach to compute a unique posterior solution of the eSIP determined by a prior distribution that encapsulates prior information about the physical system.
@@
We return to consider the eSIP when developing the estimation methodology in Sec. 3.
@@ 1.2 An overview of the main results
Assume that $P_{\mathcal{D}}$ has density $\rho_{\mathcal{D}}$ with respect to Lebesgue measure. Under general conditions, for each prior distribution $P_{\mathrm{p}}$ on $(\Lambda,\mathcal{B}_{\Lambda})$ with density $\rho_{\mathrm{p}}$ there is a unique posterior ${P}_{\Lambda}$ that solves the SIP and there is an explicit formula for its density ${\rho}_{\Lambda}$ determined from $\rho_{\mathcal{D}}$ and $\rho_{\mathrm{p}}$. This is the nonparametric analog of the result typically deployed for Bayesian statistics for a parametric statistical model.
@@
Under general conditions, the density ${\rho}_{\Lambda}$ of the posterior is continuous a.e.
@@
Under general conditions, the posterior corresponding to the uniform prior has maximum entropy.
@@
Returning to the eSIP, we construct an estimator $\hat{P}_{\Lambda}\approx P_{\Lambda}$ using a form of importance sampling. Under general conditions, the estimator $\hat{P}_{\Lambda}$ is asymptotically unbiased and strongly consistent and we provide asymptotic estimates of its variance and error.
Broadly speaking, the first result says that even for complex computer models, there exists a unique posterior with an explicit formula for the density. The second result implies that the posterior is well-behaved in practical settings. The third main result says that under general conditions that the uniform prior is the least informative choice. Finally the fourth main result is that there is a robust, reliable, and practical nonparametric estimator.
@@
We estimate the posterior corresponding to the uniform prior using the nonparametric estimator presented in Sec. 3 with a large number of samples in order to virtually eliminate effects of finite sampling. We sample $16\times 10^{6}$ points in $\Lambda$ randomly from $P^{\mathrm{d}}_{\Lambda}$. We use the synthetic data to build an empirical output distribution on a partition of $\mathcal{D}$ with $100$ cells. By a partition of a measurable set $A$, we mean a finite or countable collection of non-intersecting measurable subsets $\{B_{i}\}_{i}$ such that $A=\bigcup_{i}B_{i}$, except possibly for set of measure 0. We compute an empirical estimator of the posterior using $16\times 10^{6}$ uniformly distributed points in $\Lambda$.
@@
In Fig. 3, we plot heatmaps of the estimators for $T=.5$ and $T=2$ produced by computing $\{P_{\Lambda}(B_{i})\}_{i=1}^{6400}$ for a partition $\{B_{i}\}_{i=1}^{6400}$ of $\Lambda$ into small cells $B_{i}$ and generating a heatmap from the probabilities. Recalling the contours shown in Fig. 1, the posteriors have the characteristic “ridge” shape oriented along contours that reflects the fact that parameter values on the same contour have equal probability due to the choice of the uniform prior. This reflects the indeterminacy induced by the loss of information in the experimental observation.
@@
Figure 3: Left and center: Heatmaps of estimators of the posterior for the exponential decay model using the uniform prior for $T=.5$ (left) and $T=2$ (center). Right: Overlays of the two heatmaps. Event A is relatively low probability for both posteriors, Event B is relatively high probability for the posterior for time .5.5 but relatively low probability for the posterior for time 22, while Event C is the reverse. Event D is relatively high probability for both posteriors.
We do not expect to recover the data generating distribution in general. We estimate a posterior determined by the choice of prior. To give intuition about the connection between the data generating distribution and the posterior, we plot the samples from the data generating distribution that produced the simulated output data together with the posterior for $T=2$ in Fig. 4. Any point on the generalized contour determined by a sample from the data generating distribution yields the corresponding observed datum. Choosing the uniform prior means that the posterior “extends” the profile of the data generating distribution along the generalized contours. In Example 7, we show the use of an informed prior that results in a posterior with more localized support.
Figure 4: The posterior for the exponential decay model using the uniform prior for $T=2$ together with the sample points (red) from the data generating distribution used to create the simulated output data.
1.3 Using the solution of calibration
@@
In applications of the calibration of computer models, the principal inferential target is information about the probability of parameters consistent with the model and the observations. The posterior solution of the eSIP is a probability distribution on the sample space $\Lambda$ that is evaluated to estimate probabilities of events. For example in the right hand plot in Fig. 3, we show four events of equal size located in different parts of $\Lambda$. We see that event $A$ has relatively low probability with respect to the posteriors for $T=.5$ and $T=2$, while event $D$ has relatively high probability for both posteriors. Events $B$ and $C$ are relatively high probability for one of the posteriors but not the other.
@@
We emphasize that the posterior solution of the SIP is a probability distribution, not a point estimate. We can compute statistics of the posterior such as a mean or variance but those are typically of little use.
@@
The other use of the posterior is forecasting or prediction, which means computing a probability distribution on a new quantity of interest. For example, we can estimate the posterior conditioned on data observed at an early time, then use the posterior to estimate the forward distribution on a QoI at a later time. This is related to ensemble forecasting, see [47] for an example of forecasting infections of COVID during a surge.
@@ 1.4 Observations and related work
Bayesian statistics, classic parametric Bayesian calibration, and nonparametric Bayesian calibration, equivalently eSIP, share a common goal of estimating a distribution for a dataset. The first two adopt a parametric statistics model to estimate the distribution while the latter employs a nonparametric estimate. All three employ conditional probability conditioned on data, so they share the common theoretical foundation of regular conditional probability conditioned on a random variable [42, 8, 44, 4]. Bayesian statistics and parametric Bayesian calibration employ Bayes’ rule, which means introducing random variables for the distribution model and the prior and the corresponding notation follows [64]. Nonparametric Bayesian calibration uses disintegration of measures with a corresponding notation [42, 21, 64, 8, 44, 59, 34, 4]. Unfortunately, while both notations originate in statistics, they are not well-aligned. We have supplied a couple of simple “hand computation” examples to illustrate disintegration and its notation.
In nonparametric Bayesian calibration, the regular conditional probability resulting from conditioning on the sub-$\sigma$-algebra induced by $Q$ plays a prominent role in construction of the posterior conditioned on the data. This reflects the fact that the experimental data provides limited information about the physical parameters. Typically, informal descriptions of Bayesian statistics suppress the role of the random vectors in conditional probabilities as those details are generally not critical for the results. However, classic parametric Bayesian computer model calibration combines experimental data and values of the computer model to estimate a posterior for the parameters (“calibration parameters”) in the presence of uncertainty [45, 39, 40, 71, 58, 38, 52, 75, 7]. So, parametric and nonparametric Bayesian calibration consider similar inverse problems, see [38]. Classic Bayesian calibration often employs a statistical surrogate model for the computer model called an emulator and attempts to model the systematic “discrepancy” between the calibrated model emulator and the mean of the physical system.
@@
The data for the nonparametric Bayesian calibration results from performing trials of the experiment multiple times, with each trial conducted with physical parameters distributed according to an unknown data generating distribution. The stochastic variation in output data is the result of random sampling of the physical parameters rather than from observational noise. Nonparametric Bayesian calibration is closely related to ensemble forecasting, e.g., as used in weather prediction [27, 1, 51, 50, 49, 35, 56, 47]. In the common situation in which the sample space $\Lambda$ of physical parameters includes ranges of values that lead to widely varying behavior in the system, e.g., bifurcation points, the posterior is a probability distribution with a complex heterogeneous structure. This motivates the use of a nonparametric estimator. Point estimates and credibility intervals are often useful in Bayesian statistics, but not usually useful for the eSIP.
The production of data for the nonparametric Bayesian calibration is different than the situation in which data results from noisy observations of a single trial corresponding to a single value of the physical parameters. A stochastic approach to the latter problem widely employed in engineering [31, 24, 67, 28, 19, 69, 70] is the Bayesian Inverse Problem. This approach involves the introduction of observational noise, regularization of the physical model, and an approach to choice of prior that is generally different from that of Bayesian statistics. The review paper [6] explores the relationship between the eSIP and the Bayesian Inverse Problem. A variation arises when longitudinal data for a single trials is available, e.g., 4D-VAR, data assimilation, ensemble Kalman filtering, Bayesian particle filtering, et cetera [1, 33, 2, 23, 3, 20, 30, 30, 72, 26, 62]. These approaches also involve the introduction of random noise and regularization.
@@
The original development of the SIP used a different form of disintegration and employed a deterministic estimate methodology [9, 11, 12]. This paper introduces a new form of disintegration for the solution that enables derivation of some key properties for the first time as well as analysis of a novel estimator based on random sampling. The advances presented in this paper depend on the developments in the theses [74, 22]. Extensions of the eSIP to other problems are considered in [54, 17, 74, 22]. Consideration of computational aspects of the solution of the eSIP are considered in [18, 15, 16, 66, 77, 78]. Applications of the eSIP can be found in [10, 14, 13, 53, 36, 43, 47]. In traditional inverse problems, “ill-conditioning” and the condition number play important roles. In [66], we develop a theory of condition for the eSIP and explore the consequences for the solution. In [5], we consider the inclusion of random effects in the eSIP, e.g., noisy data, which opens up doors to numerous applications and extensions, including connecting the eSIP more closely to traditional Bayesian statistics.